The thermodynamic solubility product of AgCl is 1.6 x 10-10. What is [Ag+] in (a) a 0.020 molar KNO3 solution and (b) a 0.020 molar KCl solution?

Need help on how to solve this problem, Thanks!!
Answer #1

a) 1.6 x 10-10 = [Ag+][Cl-]

The presence of KNO3 won't change the concentration of any of them, so we will proceed as follows:

1.6 x 10-10 = x2

x = 1.2649 x 10-5 M = [Ag+]

For case b)

1.6 x 10-10 = [Ag+] [0.02]

[Ag+] = 8 x 10-9